contest consists of finding all of the code words that can be formed from the letters in the name "ATARI." Assume that the letter A can be used twice, but the others at most once.
A) How many five-letter words can be formed?
B) How many two-letter words can be formed?
C) How many words can be formed?

Found 3 solutions by math_tutor2020, ikleyn, greenestamps:
Answer by math_tutor2020(3817)


Part (a)

If we could tell the "A"s apart, then there would be 5! = 5*4*3*2*1 = 120 different five-letter permutations.

But because we cannot tell the "A"s apart, we must divide by 2 which corrects the erroneous double-count.

We'll have 120/2 = 60 different five-letter words possible.

-------------------------------

Part (b)

We have these four possible cases:
  1. The letter "A" is used twice
  2. The letter "A" is used exactly once in the 1st slot.
  3. The letter "A" is used exactly once in the 2nd slot.
  4. The letter "A" is not chosen at all.
Case 1 has exactly one outcome. That outcome is AA.

Case 2 has 3 choices for the 2nd slot, which means we have the 3 permutations AT, AR, AI

Case 3 also has 3 outcomes. The logic is similar to case 2.

Case 4 has 3*2 = 6 permutations of the letters {T,R,I}

Add up the results
1+3+3+6 = 13
There are 13 different two letter sequences


The entire list is
  1. AA
  2. AT
  3. AR
  4. AI
  5. TA
  6. TR
  7. TI
  8. RA
  9. RT
  10. RI
  11. IA
  12. IT
  13. IR
A combinatorics calculator I recommend is this
https://www.mathsisfun.com/combinatorics/combinations-permutations-calculator.html
It computes the nPr value, and also lists the different permutations.
Refer to the "List Them" section on that page.

-------------------------------

Part (c)

The previous part calculated 13 different two-letter words.

Let's calculate how many three-letter sequences are possible.

We have these cases
  1. The letter "A" is used twice.
  2. The letter "A" is used exactly once.
  3. The letter "A" is not chosen at all.


Case 1
There are 3 slots to pick from where the letters T, R, or I will go.
After that slot is chosen, the other two slots are locked in as "A"
Therefore, we have 3*3 = 9 ways to do case 1.

Case 2
There are 3 slots to pick for the letter "A".
We have two more slots and 3 more letters to pick from to fill said slots. We'll have 3*2 = 6 permutations of those other letters.
Overall we have 3*6 = 18 ways to carry out case 2.

Case 3
There are 3*2*1 = 6 permutations of the letters {T,R,I}

Add up the results
9+18+6 = 33
We find there are 33 different three-letter words.

I'll let you do the scratch work to determine how many four-letter words are possible.
You should find there are 60 different four-letter words possible.

To summarize we have...
  • 60 ways to form a five-letter word.
  • 60 ways to form a four-letter word.
  • 33 ways to form a three-letter word.
  • 13 ways to form a two-letter word.
There are 60+60+33+13 = 166 different words possible when we can only select letters from the list {A,T,A,R,I} where "A" can be used at most twice, but the other letters at most once.

-------------------------------


Answers:
(a) 60
(b) 13
(c) 166

Answer by ikleyn(52787)

.

I think that the answer  166  to question  (c)  in the post by  @math_tutor2020 should be changed:

4  other  1-letter words should be added to it

            166 + 4 = 170.



Answer by greenestamps(13200)



1-letter words....
There are 4 different letters to choose from.
The number of 1-letter words is 4.

2-letter words....
If both letters A are used, there is only 1 word -- AA
If the letter A is not repeated, then there are 4 choices for the first letter and then 3 choices for the second, so there are 4*3=12 2-letter words without both letters A.
The number of 2-letter words is 1+12 =13.

3-letter words....
If both letters A are used, then the combination of letters can be AAI, AAR, or AAT. There are (3!)/(2!) = 3 arrangements of each of those, for a total of 9.
If both letters A are not used, then there are 4 choices for the first letter, 3 for the second, and 2 for the third, for a total of 4*3*2 = 24.
The number of 3-letter words is 9+24 = 33.

4-letter words....
If both letters A are used, then the combination of letters can be AAIR, AAIT, or AART. There are (4!)/(2!) = 12 arrangements of each of those, for a total of 36.
If both letters A are not used, then the letters are ATRI in any order; there are 4! = 24 of those.
The number of 4-letter words is 36+24 = 60.

5-letter words....
The letters are ATARI in any order, the number of arrangements is (5!)/(2!) = 60.
There are 60 5-letter words.

Summary:
1-letter words: 4
2-letter words: 13
3-letter words: 33
4-letter words: 60
5-letter words: 60

Total number of words: 4+13+33+60+60 = 170

ANSWERS:
A) 5-letter words: 60
B) 2-letter words: 13
C) total number of words: 170